# Control of the local and non-local electromagnetic response in all-dielectric reconfigurable metasurfaces

### v1

@@ II Electric and magnetic response tunability
Figure 2 (d) shows the transmittance, defined as the absolute value of the transmission coefficient squared. The colormap depicts the dependency on the height of the nanobars when the period of the structure is fixed to $D=0.52\;\lambda_{\rm W}$, the width $w=0.27\;\lambda_{\rm W}$ and the refractive index of the PCM is $n_{\rm PCM}=n_{\rm Ge}=4$. Notice that in this configuration, $n_{\rm PCM}=n_{\rm Ge}$ and the structure can be modeled as a homogeneous nanobar made of germanium. The resonant modes produce zero transmission, and their dependency on the parameter $h$ is marked with white dashed lines. As expected, an increment in the height produces a displacement of the three resonant modes towards higher wavelengths. From this analysis, one can see that the two electric modes, $\mathrm{ED}_{x}$ and $\mathrm{ED}_{z}$, change differently with the parameter $h$, allowing the overlap of both resonances at the same frequency. As it was demonstrated, in [29], zero backscattering and full control of the phase in transmission could be obtained by bringing together the symmetric (respect to the $z$-axis) $\mathrm{ED}_{x}$ resonance and the antisymmetric $\mathrm{ED}_{z}$. Under this condition, the structure forms a Huygens pair that can be used to control the direction of propagation of the transmitted EM waves.
@@
Once the resonance pair is constructed, the geometrical parameters and $\lambda_{\rm W}$ are left fixed, and we can study the effect of changing the refractive index in the PCM layer. The height of the PCM is set to be $h_{\rm PCM}=h/3$ and $p=h/3$. In Fig. 3, the phase and transmittance of the metasurface are represented for the range of $n_{\rm PCM}$ between 3 and 6.5. Transmittance is near unity for all the refractive index values, Fig. 3(b). Beyond the limit in refractive index ($n_{\rm PCM}>6.5$), the PCM layer is so electrically large that the displacement field produces a resonance inside this layer, detuning the Huygens’ pair. When ($6.53<n_{\rm PCM}<$ 6.5), the change of the refractive index in the central slab shifts each ED resonance by roughly the same amount so the pair is close enough to create high transmittance while changing the phase drastically. The phase coverage in the represented range is almost complete, Fig. 3(a). It is important to notice that the chosen configuration, a central layer of PCM sandwiched between two dielectric germanium layers, is symmetric in the $z$-direction, neglecting EM coupling and maintaining high transmission. The resonance pair is stronger in the center of the rectangular bar; therefore, the refractive index change has more effect on both ED resonances and can achieve a great change in the phase of the transmitted wave. Furthermore, the geometry prevents polarization conversion, as shown in Ref. [29].
@@ III Electromagnetic coupling tunability
In the case of the structure shown in Fig. 6, the response can be modeled by using only a representation in cartesian components of the tangential part of polarizability tensors, as the metasurface is normally illuminated and the induced moments are going to be restricted to the $x-y$ plane. Furthermore, the geometry presented is non-chiral and is made out of an isotropic material, and the lattice only presents $\rm C_{2}$ symmetry, ensuring no cross-polarization conversion on a linear basis. In the following, the notation used is going to have the form $t_{ij}^{k}$ for transmission parameters or $r_{ij}^{k}$ for reflection parameters, where superscripts $i,j=\{+,-\}$ indicate the direction of propagation $+z$ or $-z$, respectively, and subscripts $k=\{x,y\}$ refer to the polarization of the incoming and outcoming field. Due to the discussed symmetry restrictions, the scattering matrix has some of its elements vanishing, $t_{xy}^{\pm}=t_{yx}^{\mp}=0$ and $r_{xy}^{\pm}=r_{yx}^{\pm}=0$. Moreover, due to reciprocity, $t_{xx}^{+}=t_{xx}^{-}$ and $t_{yy}^{+}=t_{yy}^{-}$. Therefore, only six parameters are independent in the scattering response problem: $t_{xx}$, $t_{yy}$, $r_{xx}^{+}$, $r_{xx}^{-}$, $r_{yy}^{+}$, and $r_{yy}^{-}$. Notice how, for the transmission coefficients, we have dropped the superscripts as reciprocity makes the direction of propagation irrelevant.
@@
The results for the absorption of the metasurface, when illuminated by x-polarized plane waves at normal incidence, are shown in Fig. 7. The crystallization, $c$, state is represented with different colors to show how, by tuning this magnitude, using heating devices or optical pumping, the absorption from one side of the metasurface can be changed from $9.7\%$ to $98.6\%$ at 193 THz while for illumination from the opposite side, it changes from $15.2\%$ to $31.0\%$. Therefore, at that frequency, the value of the contrast between each side’s absorptions reaches $\Delta A_{x}=-0.676$ for $c=1$. The parameter $c$ ranges from 0 to 1 and accounts for the fraction of material that has been crystallized.
@@ IV Conclusions
By harnessing the principles of a Huygens’ source, this configuration offers practical applications in designing metasurfaces that can alter the direction of transmitted waves. This is achieved through a local approach, implementing a linear phase gradient along the metasurface. The ability to control the number of resonators used to create the gradient allows for beam steering. Moreover, the metasurface can be programmed to utilize auxiliary evanescent fields, creating strong non-local effects that can tailor the energy to different diffraction modes, thereby surpassing the limitations of local designs.

### v2

@@ II Electric and magnetic response tunability
Figure 2 (d) shows the transmittance, defined as the absolute value of the transmission coefficient squared. The colormap depicts the dependency on the height of the nanobars when the period of the structure is fixed to $D=0.52\;\lambda_{\rm W}$, the width $w=0.27\;\lambda_{\rm W}$ and the refractive index of the PCM is $n_{\rm PCM}=n_{\rm Ge}=4$. Notice that in this configuration, $n_{\rm PCM}=n_{\rm Ge}$ and the structure can be modeled as a homogeneous nanobar made of germanium. The resonant modes produce zero transmission, and their dependency on the parameter $h$ is marked with white dashed lines. As expected, an increment in the height produces a displacement of the three resonant modes towards higher wavelengths. From this analysis, one can see that the two electric modes, $\mathrm{ED}_{x}$ and $\mathrm{ED}_{z}$, change differently with the parameter $h$, allowing the overlap of both resonances at the same frequency. As it was demonstrated, in [29], zero backscattering and full control of the phase in transmission could be obtained by bringing together the symmetric (respect to the $z$-axis) $\mathrm{ED}_{x}$ resonance and the antisymmetric $\mathrm{ED}_{z}$. Under this condition, the structure forms a Huygens’ pair that can be used to control the direction of propagation of the transmitted EM waves.
@@
Once the resonance pair is constructed, the geometrical parameters and $\lambda_{\rm W}$ are left fixed, and we can study the effect of changing the refractive index in the PCM layer. The height of the PCM is set to be $h_{\rm PCM}=h/3$ and $p=h/3$. In Fig. 3, the phase and transmittance of the metasurface are represented for the range of $n_{\rm PCM}$ between 3 and 6.5. Transmittance is near unity for all the refractive index values, Fig. 3(b). Beyond the limit in refractive index ($n_{\rm PCM}>6.5$), the PCM layer is so electrically large that the displacement field produces a resonance inside this layer, detuning the Huygens pair. When ($6.53<n_{\rm PCM}<$ 6.5), the change of the refractive index in the central slab shifts each ED resonance by roughly the same amount so the pair is close enough to create high transmittance while changing the phase drastically. The phase coverage in the represented range is almost complete, Fig. 3(a). It is important to notice that the chosen configuration, a central layer of PCM sandwiched between two dielectric germanium layers, is symmetric in the $z$-direction, neglecting EM coupling and maintaining high transmission. The resonance pair is stronger in the center of the rectangular bar; therefore, the refractive index change has more effect on both ED resonances and can achieve a great change in the phase of the transmitted wave. Furthermore, the geometry prevents polarization conversion, as shown in Ref. [29].
@@ III Electromagnetic coupling tunability
In the case of the structure shown in Fig. 6, the response can be modeled by using only a representation in cartesian components of the tangential part of polarizability tensors, as the metasurface is normally illuminated and the induced moments are going to be restricted to the $x$-$y$ plane. Furthermore, the geometry presented is non-chiral and is made out of an isotropic material, and the lattice only presents $\rm C_{2}$ symmetry, ensuring no cross-polarization conversion on a linear basis. In the following, the notation used is going to have the form $t_{ij}^{k}$ for transmission parameters or $r_{ij}^{k}$ for reflection parameters, where subscripts $i,j=\{+,-\}$ indicate the direction of propagation $+z$ or $-z$, respectively, and superscripts $k=\{x,y\}$ refer to the polarization of the incoming and outcoming field. Due to the discussed symmetry restrictions, the scattering matrix has some of its elements vanishing, $t_{xy}^{\pm}=t_{yx}^{\mp}=0$ and $r_{xy}^{\pm}=r_{yx}^{\pm}=0$. Moreover, due to reciprocity, $t_{xx}^{+}=t_{xx}^{-}$ and $t_{yy}^{+}=t_{yy}^{-}$. Therefore, only six parameters are independent in the scattering response problem: $t_{xx}$, $t_{yy}$, $r_{xx}^{+}$, $r_{xx}^{-}$, $r_{yy}^{+}$, and $r_{yy}^{-}$. Notice how, for the transmission coefficients, we have dropped the superscripts as reciprocity makes the direction of propagation irrelevant.
@@
The results for the absorption of the metasurface, when illuminated by $x$-polarized plane waves at normal incidence, are shown in Fig. 7. The crystallization state, $c$, is represented with different colors to show how, by tuning this magnitude, using heating devices or optical pumping, the absorption from one side of the metasurface can be changed from $9.7\%$ to $98.6\%$ at 193 THz while for illumination from the opposite side, it changes from $15.2\%$ to $31.0\%$. Therefore, at that frequency, the value of the contrast between each side’s absorptions reaches $\Delta A_{x}=-0.676$ for $c=1$. The parameter $c$ ranges from 0 to 1 and accounts for the fraction of material that has been crystallized.
@@ IV Conclusions
By harnessing the principles of a Huygens source, this configuration offers practical applications in designing metasurfaces that can alter the direction of transmitted waves. This is achieved through a local approach, implementing a linear phase gradient along the metasurface. The ability to control the number of resonators used to create the gradient allows for beam steering. Moreover, the metasurface can be programmed to utilize auxiliary evanescent fields, creating strong non-local effects that can tailor the energy to different diffraction modes, thereby surpassing the limitations of local designs.